Can anyone help with this problem with detail solutions? Thank you!

The problem is that you need to borrow money to start the business and your banker has asked for a break- even analysis. YouPart 1: 1. Whats the break even point in units--how many dogs will you need to groom in the first year to break even? 2. If
Answer #1

Variable Costs : 6.00 $

Fixed Costs: 146000

Part 1:

1. Contribution per unit = 32 - 6 = 26, BEP in units = 146000/26 = 5615 dogs

2. Total available hours: 3 person * 7 hours * 5 days * 50 weeks = 5250 hours, So, 5615/5250 = 1.06 dogs per hour. Each have 7 hours in a day. So, 7.5 (or 8) dogs. It requires more than a Dog in an hour.

Part 2:

1. Contribution per unit = 38 - 6 = 32, BEP in units = 146000/32 = 4563 dogs

2. So, 4563/5250*7 hours = 6.08 (or 6) Dogs

Part 3:

1. Yes I would start this business.

2. Price depends should be such that it recovers the variable costs first. and then recover the fixed costs. the practical prices range from 30$ to 90$.

3. Break even point can be reduced by reducing the fixed costs or by reducing the variable costs per unit or by increasing the grooming fee.

Please comment in case of any query.